OpenStudy (anonymous):

Sharita does typing jobs for scientists, She charges $10 for the first 3 pages & $3 for each additional page,Sonia paid $97 to have a research paper typed, how many pages long was the paper?

hero (hero):

Hint: 3x + 30 = 97

hero (hero):

Kinda tricky because when you find x, you still have to add the extra three pages

OpenStudy (anonymous):

x=10?

OpenStudy (anonymous):

@Hero

hero (hero):

3x + 30 = 97 3x = 67 x = 67/3 x = 23 and 1/3 Plus the 3 pages makes 26 and 1/3 pages

hero (hero):

So the answer is 26 and 1/3 pages